# Data-adaptive structural change-point detection via isolation

### v1

@@ Appendix D Proof of Theorem 2
where $C^{\ast}$ satisfies assumption (A3), $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{logT}$ and $3/2\leq n\leq\delta_{T}/2\lambda_{T}$. As before, for $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). For $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\}$ being the location of the largest difference detected in the interval $[s,e]$, $1\leq s<e\leq T$, define $c_{m}^{l}$ and $c_{k}^{r}$ as in (B). Since the length of the intervals in (5) is $\delta_{T}/2n$ and $\lambda_{T}\leq\delta_{T}/2n$, for $m,k\in\{0,1,\ldots,K^{\max}\}$ we ensure that there exists at least one $m$ and at least one $k$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$.
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point, where $\delta_{1,T}^{j}\leq\frac{\delta_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\delta_{j}=|r_{j+1}-r_{j}|$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3.
Without loss of generality, we suppose that the first change-point to get detected is $r_{J}$ for some $J\in\left\{1,2,\ldots,N\right\}$. We can show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that $r_{J}$ is isolated using exactly the same argument as in (B). We will now show that for $\tilde{b}_{J}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C^{t}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})$, then $C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})>\zeta_{T}$. Using (8), we have that
@@
Now, concentrating on the case that the next change-point to get detected is $r_{J-1}$, we mention that since this is the closest change-point to the already detected $r_{J}$, we need to make sure that detection is possible. As before, for $k_{J-1},m_{J-1}\in\{0,1,\ldots,K^{\max}\}$ we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$ and its detection is at location
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (B). Using Lemma 3, we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}C^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\left(\boldsymbol{X}\right)$,

### v2

@@ Appendix D Proof of Theorem 2
where $C^{\ast}$ satisfies assumption (A3), $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{\log T}$ and $3/2\leq n\leq\delta_{T}/2\lambda_{T}$. As before, for $j\in\{1,2,\ldots,N\}$ define $I^{L}_{j}$ and $I^{R}_{j}$ as in (5). For $d_{s,e}=\textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\{\lvert X_{t+2}-2X_{t+1}+X_{t}\rvert\}$ being the location of the largest difference detected in the interval $[s,e]$, $1\leq s<e\leq T$, define $c_{m}^{l}$ and $c_{k}^{r}$ as in (C). Since the length of the intervals in (5) is $\delta_{T}/2n$ and $\lambda_{T}\leq\delta_{T}/2n$, for $m,k\in\{0,1,\ldots,K^{\max}\}$ we ensure that there exists at least one $m$ and at least one $k$ such that $c_{m}^{l}\in I^{L}_{j}$ and $c_{k}^{r}\in I^{R}_{j}$ for all $j\in\{1,2,\ldots,N\}$.
At the beginning of DAIS, $s=1,e=T$ and the first change-point that will get detected depends on the value of $d_{1,T}$. As already explained in Section 2.3, for $j\in\{1,\ldots,N-1\}$, the largest difference $d_{s,e}$ will be at most at a distance $\delta_{1,T}^{j}$ from the nearest change-point $r_{j}$ or $r_{j+1}$, where $\delta_{1,T}^{j}\leq\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}$ for $\tilde{\delta}_{j}=r_{j+1}-r_{j}$. The first point to get detected will be the point that is closest to the largest difference $d_{1,T}$. The interval where the detection of this change-point occurs, cannot contain more than one change-points, as was explained in Section 2.3.
Without loss of generality, we suppose that the first change-point to get detected is $r_{J}$ for some $J\in\left\{1,2,\ldots,N\right\}$. We can show that there exists an interval $[c_{m}^{l},c_{k}^{r}]$, for $m,k\in\{0,1,\ldots,K^{\max}\}$, such that $r_{J}$ is isolated using exactly the same argument as in (C). We will now show that for $\tilde{b}_{J}=\textrm{argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C^{t}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})$, then $C^{\tilde{b}_{J}}_{c_{m}^{l},c_{k}^{r}}(\boldsymbol{X})>\zeta_{T}$. Using (D), we have that
@@
Now, concentrating on the case that the next change-point to get detected is $r_{J-1}$, we mention that since this is the closest change-point to the already detected $r_{J}$, we need to make sure that detection is possible. As before, for $k_{J-1}\in\{1,\ldots,K^{\max}\}$ and $m_{J-1}\in\{k_{J-1}-1,k_{J-1}\}$ we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$ and its detection is at location
@@
which satisfies $\bigl{|}r_{J-1}-b_{J-1}\bigl{|}\Bigl{(}\Delta_{J-1}^{f}\Bigl{)}^{2/3}\leq C_{3}\left(\log T\right)^{1/3}$. Firstly, $r_{J-1}$ is isolated in the interval $[c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}]$ using the same argument as in (C). Using Lemma 3, we have that for $\tilde{b}_{J-1}=\textrm{argmax}_{c_{m_{J-1}}^{l}\leq t<c_{k_{J-1}}^{r}}C^{t}_{c_{m_{J-1}}^{l},c_{k_{J-1}}^{r}}\left(\boldsymbol{X}\right)$,
@@
Appendix E Outline of the proof of Theorem 3
[…]
Step 1: We show that the values of the mean-dominant norm of the contrast function of the observed and unobserved signals, $L\left(\boldsymbol{\tilde{X}_{s,e}^{b}}\right)$ and $L\left(\boldsymbol{\tilde{f}_{s,e}^{b}}\right)$, are uniformly close, for all $1\leq s\leq b<e\leq T$. Using the mean-dominance property as given in p.190 of [9], it holds that $0\leq L_{1}(\boldsymbol{x})\leq L(\boldsymbol{x})\leq L_{\infty}(\boldsymbol{x}),\forall\boldsymbol{x}\in(\mathbb{R}^{d})^{+}$ and so, for $A_{T}^{*}=\bigg{\{}\max_{1\leq s\leq b<e\leq T}L\left(\left|\boldsymbol{\tilde{X}_{s,e}^{b}}-\boldsymbol{\tilde{f}_{s,e}^{b}}\right|\right)\leq\sqrt{8\operatorname{log}\left(Td^{\frac{1}{4}}\right)}\bigg{\}}$ and $A_{T}=\bigg{\{}\max_{1\leq s\leq b<e\leq T}\left|C^{b}_{s,e}-D_{s,e}^{b}\right|\leq\sqrt{8\operatorname{log}\left(Td^{\frac{1}{4}}\right)}\bigg{\}}$, it holds that $A^{\ast}_{T}\subseteq A_{T}$. From the definition of our model in (20), we have that $\boldsymbol{\epsilon_{t}}\sim\mathcal{N}_{d}(\boldsymbol{0},\Sigma),\forall t\in\left\{1,\ldots,T\right\}$ and, using the Bonferroni inequality, we can show, following similar steps as those in [3], that $\mathbb{P}\left(\left(A^{\ast}_{T}\right)\right)\leq 1/(12\sqrt{\pi}T)$, which implies $\mathbb{P}\left(A_{T}\right)\geq\mathbb{P}\left(A^{\ast}_{T}\right)\geq 1/(12\sqrt{\pi}T)$.
Step 2: We control the distance between $L\left(\boldsymbol{\tilde{X}_{s,e}^{b_{1}}}\right)-L\left(\boldsymbol{\tilde{X}_{s,e}^{b_{2}}}\right)$ and $L\left(\boldsymbol{\tilde{f}_{s,e}^{b_{1}}}\right)-L\left(\boldsymbol{\tilde{f}_{s,e}^{b_{2}}}\right)$, for all possible combinations of $s,e,b_{1},b_{2}$, where $1\leq s<e\leq T$. For $\boldsymbol{f_{k}}=\left(f_{1,k},\ldots,f_{T,k}\right)$ and for $[s,e]$ being any interval that contains only one true change-point, namely $r_{j}$, we denote
$A_{s,e}^{b}(k,r_{j}):=\frac{\left\langle\boldsymbol{\psi_{s,e}^{b}}\left\langle\boldsymbol{f_{k}},\boldsymbol{\psi_{s,e}^{b}}\right\rangle-\boldsymbol{\psi_{s,e}^{r_{j}}}\left\langle\boldsymbol{f_{k}},\boldsymbol{\psi_{s,e}^{r_{j}}}\right\rangle,\boldsymbol{\epsilon}\right\rangle}{\left\|\boldsymbol{\psi_{s,e}^{b}}\left\langle\boldsymbol{f_{k}},\boldsymbol{\psi_{s,e}^{b}}\right\rangle-\boldsymbol{\psi_{s,e}^{r_{j}}}\left\langle\boldsymbol{f_{k},\boldsymbol{\psi_{s,e}^{r_{j}}}}\right\rangle\right\|_{2}},$ (62)
where $\psi_{s,e}^{b}$ is defined in (27). In (62) above, $k\in\left\{1,\ldots,d\right\}$, while $\|\cdot\|_{2}$ is the Euclidean norm. Due to $\epsilon_{t,k},t=1,\ldots,T$ following the standard normal distribution for all $k\in\left\{1,\ldots,d\right\}$ (with the univariate error terms being independent over time, but possibly spatially dependent), then straightforward calculations lead to $A_{s,e}^{b}(k,r_{j})\sim\mathcal{N}(0,1),\forall k\in\left\{1,\ldots,d\right\}$. For
$B_{T}=\left\{\max_{\begin{subarray}{c}j=1,\ldots,N\\ k=1,\ldots,d\end{subarray}}\max_{\begin{subarray}{c}r_{j-1}<s\leq r_{j}\\ r_{j}<e\leq r_{j+1}\\ s\leq b<e\end{subarray}}\left|A_{s,e}^{b}(k,r_{j})\right|\leq\sqrt{8\operatorname{log}\left(Td^{\frac{1}{4}}\right)}\right\},$ (63)
we can show that $\mathbb{P}\left(B_{T}\right)\geq 1-1/(12\sqrt{\pi}T)$ using a similar procedure as in Step 1. From Steps 1 and 2, we conclude that $\mathbb{P}\left(A_{T}^{*}\cap B_{T}\right)\geq 1-\frac{1}{6\sqrt{\pi}T}.$
Step 3: Working under the assumption that both $A^{\ast}_{T}$ (and thus also $A_{T}$) and $B_{T}$ hold, we show that detection in an interval where the change-point is isolated will occur and so the proof can be restricted to an interval with a single change-point and we also prove that the estimated and true change-points are close to each other. The values used in this step for the positive constants $C_{1},C_{2},$ and $C_{3}$ of Theorem 3, are $C_{1}=\sqrt{C_{3}}+\sqrt{8},C_{2}=\frac{1}{\sqrt{4n}}-\frac{2\sqrt{2}}{\underline{C}_{M}},C_{3}=2(2\sqrt{2}+4)^{2}$. The order with which the change-points are detected depends on the location of the largest difference at each step of the algorithm and so, without loss of generality, we assume that the first change-point to get detected is $r_{J}$ for $J\in\{1,\ldots,N\}$. For $c_{m}^{l},c_{k}^{r}$ as defined in (C), $c_{m}^{l}\in I_{J}^{L}$ or $c_{k}^{r}\in I_{J}^{R}$ hold (or both), with $I_{J}^{L},I_{J}^{R}$ as defined in (5), and showing that the change-point will be isolated in $[c_{m}^{l},c_{k}^{r}]$ is the same as for Theorem1 Step 3. It can be shown that for $\tilde{b}={\rm argmax}_{c_{m}^{l}\leq t<c_{k}^{r}}C_{c_{m}^{l},c_{k}^{r}}^{t}$, $C_{c_{m}^{l},c_{k}^{r}}^{\tilde{b}}>\zeta_{T,d}$, so $r_{J}$ can be detected in the interval.
Now, for $k^{\ast}\in\{1,\ldots,K^{\max}\}$ and $m^{\ast}\in\{k^{\ast}-1,k^{\ast}\}$, we denote by $c_{m^{\ast}}^{l}\geq c_{m}^{l}$ and $c_{k^{\ast}}^{r}\leq c_{k}^{r}$ the first left- and right-expanding points, respectively, where detection of $r_{J}$ occurs. Showing that the estimated and true change-points are close to each other requires working on the two mean-dominant norms of (21) separately. Denoting by $q_{J}:={\rm argmax}_{j=1,\ldots,d}\left|\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b_{J},j}\right|$ and $\alpha=\begin{cases}\Delta^{q_{J}}_{J},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ L_{2}(\boldsymbol{\Delta_{J}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases}$, we show that $\left|b_{J}-r_{J}\right|\alpha^{2}\leq C_{3}\log\left(Td^{\frac{1}{4}}\right).$ This can be done using a contradiction argument, based on the fact that for any $b^{*}\in\left\{c_{m^{\ast}}^{l},\ldots,c_{k^{\ast}}^{r}-1\right\}$ with $\left|b^{*}-r_{J}\right|\alpha^{2}>\log\left(Td^{\frac{1}{4}}\right)$, it holds that $\left(\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{r_{J},q_{J}}\right)^{2}>\left(\tilde{X}_{c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}}^{b^{*},q_{J}}\right)^{2}$, which, using the definition of $b_{J}$, cannot be true. The detailed proof of this is skipped, but we highlight that it basically relies on a combination of the strategy followed in [3] that the consistency of the MID algorithm was proved, and the approach followed in our Theorem 1, Step 3.
Step 4: We show that when the algorithm re-starts, no change-points are left undetected and that the new intervals allow for the detection of the remaining change-points. The first part is immediate, as from Step 3 we know that the change-point was isolated in the interval in which detection occurred. Since the change-point was detected in the interval $[c_{m^{\ast}}^{l},c_{k^{\ast}}^{r}]$, MDAIS restarts on $[1,c_{m^{\ast}}^{l}]$ and $[c_{k^{\ast}}^{r},T]$. The proof that detection of $r_{j}$ occurs is the same as in Step 3 for all $j\in\{1,\ldots,J-2,J+2,\ldots,N\}$. We focus on $j=J-1$ and similar steps hold for $j=J+1$. For $k_{J-1}\in\{1,\ldots,K^{\max}\}$ and $m_{J-1}\in\{k_{J-1}-1,k_{J-1}\}$, we will show that $r_{J-1}$ gets detected in $[c_{m_{J-1}^{\ast}}^{l},c_{k_{J-1}^{\ast}}^{r}]$, where $c_{m_{J-1}^{\ast}}^{l}\geq c_{m_{J-1}}^{l}$ and $c_{k_{J-1}^{\ast}}^{r}\leq c_{k_{J-1}}^{r}\leq c_{m^{\ast}}^{l}$. The upper bound of the last inequality is the reason why $r_{J-1}$ should be considered separately, as the right end-point of the interval where detection will occur is upper bounded by the left end-point of the interval in which $r_{J}$ was detected. The detection of $r_{J-1}$ is at $b_{J-1}={\rm argmax}_{c_{m^{\ast}_{J-1}}^{l}\leq t<c_{k^{\ast}_{J-1}}^{r}}C_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{t}$, which, for $q_{J-1}:={\rm argmax}_{j=1,2,\ldots,d}\left|\tilde{X}_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{b_{J-1},j}\right|$, satisfies
$C_{3}\log\left(Td^{1/4}\right)\geq\begin{cases}\left|b_{J-1}-r_{J-1}\right|\left(\Delta_{J-1}^{q_{J-1}}\right)^{2},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ \left|b_{J-1}-r_{J-1}\right|L_{2}^{2}(\boldsymbol{\Delta_{J-1}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases}.$
Following similar steps as in Step 3, we have that $C_{c_{m^{\ast}_{J-1}}^{l},c_{k^{\ast}_{J-1}}^{r}}^{b_{J-1}}>\zeta_{T,d}$. Now, with
$\alpha_{J-2}:=\begin{cases}\Delta_{J-1}^{q_{J-1}},\;{\rm when}\;L(\cdot)=L_{\infty}(\cdot)\\ L_{2}(\boldsymbol{\Delta_{J-1}}),\;{\rm when}\;L(\cdot)=L_{2}(\cdot)\end{cases},$
it is straightforward to show, following the same process as Step 3 of the proof, that $\left|b_{J-2}-r_{J-2}\right|\alpha_{J-2}^{2}\leq C_{3}\log\left(Td^{1/4}\right)$.
Step 5: Finally, when no change-points are left, the algorithm terminates. After detecting all the change-points at locations $b_{1},b_{2},\ldots,b_{N}$ using the intervals $[c_{m_{j}^{\ast}}^{l},c_{k_{j}^{\ast}}^{r}]$ for $j\in\{1,\ldots,N\}$, the algorithm will check all intervals of the form $[c_{k_{j-1}^{\ast}}^{r},c_{m_{j}^{\ast}}^{l}]$ and $[c_{k_{j}^{\ast}}^{r},c_{m_{j+1}^{\ast}}^{l}]$, with $c_{m_{0}^{\ast}}^{l}=1$ and $c_{k_{N+1}^{\ast}}^{r}=T$. At most $N+1$ intervals of this form, containing no change-points will be checked. Denoting by $[s^{\ast},e^{\ast}]$ any of those intervals, for $b\in\{s^{\ast},s^{\ast}+1,\ldots,e^{\ast}-1\}$,
$C_{s^{*},e^{*}}^{b}\leq D_{s^{*},e^{*}}^{b}+\sqrt{8\log\left(Td^{\frac{1}{4}}\right)}=\sqrt{8\log\left(Td^{\frac{1}{4}}\right)}<C_{1}\sqrt{\log\left(Td^{\frac{1}{4}}\right)}\leq\zeta_{T,d}$
and the algorithm terminates.